A 3 kilogram mass is travelling at 15 meters per second when friction begins to decelerate it to stop in 5 seconds. What is the magnitude of the force of friction?
1
Expert's answer
2020-11-16T07:44:19-0500

Answer

Deceleration can be obtain as

"v=u-at\\0=15-a\times5\\a=3m/s^2"

So friction force can be given

"f=ma\\f=3\times3\\f=9N"